OpenStudy (anonymous):

Which of the following is not a zero of f(x) = (x2 - 4x + 4) (x2 + 3x - 10) = 0 a) -2 b) -5 c) 2

OpenStudy (johnweldon1993):

lets plug them in ((2)² - 4(2) + 4) ((2)² + 3(2) - 10) 4 - 4 + 4 = 4 times 4 + 6 - 10 10 - 10 = 0 right? so 4 times 0 is...always 0 :)

OpenStudy (anonymous):

Thank you can you continue to help me pleeeaaassee ? I have more questions ill give you a medal and ill give my imput.

OpenStudy (johnweldon1993):

no problem :)

OpenStudy (anonymous):

(f)x= x^3-3x^2+4=0 ... For the equation given 2 has multiplicity of a) 1 b) 2 c) 3 My answer was 3 ...